Grade 10Discuss with Askiitians Tutors

If A is a point(1,2) and B is the point (2,5) find the coordinates of the point C which lies on AB produced and such that AC=2 BC

THEBE DONALD MAIKETSO
8 Years agoGrade 10

2 Answers

Bir
8 Years ago
By section formula The Coordinates of point C will be, {[(mx2+nx1)/(m+n)],[(my2+ny1)/(m+n)]}. ....A(1,2)....B(2,5). m ratio n is 2:1 ............. x1=1and x2=2.......y1=2and y2=5........ By putting this in the section formula we get
Bir
8 Years ago
By section formula The Coordinates of point C will be. X={[(mx2+nx1)/(m+n)],.Y=[(my2+ny1)/(m+n)]}. ............A(1,2)....B(2,5). m ratio n is 2:1 ............. x1=1and x2=2.......y1=2and y2=5........ By putting this in the section formula we get X=5÷3. Y=4....Hope you like the answer... You can also do this by distance formula but it takes much more time than this.....thank u